OpenStudy (anonymous):

The molar heat of combustion of ethanol is 1360kJ/mol. What will be the final temperature if 250g of water at 20 Celsius, is heated by the combustion of 2.0g of ethanol if there is a 30% heat loss to the surroundings? (The specific heat capacity of water is 4.2 J K^-1 g^-1).

OpenStudy (zpupster):

number of moles of ethanol used = 2/(2x12+6+16) = 0.0357 moles therefore find the heat released by burning this much, 0.0357x1360 = 48.571 KJ (<--- note this is in kilojolues) now 30% of this is lost , actual heat used is 70% of the above amount, 0.70x48.571= 34KJ using the equation enthalpy change = mc(change in temperature) where the enthalpy change is 34 000 J (notice you must change from KJ to J) therefore 34 000 = 250x4.2x(X-20) X-20= 32.38 X=52.38 degress celcius the final temperature is therefore 52 degress celcius

OpenStudy (anonymous):

Thankyou so much, much appreciated. But you made a little mistake with the molar mass of ethanol calculation, its supposed to be (12.01*2+1.008*6+16.00) but I understood what you meant. Thankyou very much sir!